OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 4x - y - 3z = -3 x - 2y - 2z = -3 x + y + 3z = -2

zepdrix (zepdrix):

Substitution for 3 equations?? Hmm interesting :o Ok there are many different ways we could do this. let's try this. Using the third equation, let's solve for y.\[\large x+y+3z=-2 \qquad \qquad \rightarrow \qquad \qquad \color{royalblue}{y=-2-x-3z}\] Then let's plug our new `y` into one of the other equations. How bout the first one?\[\large 4x-\color{royalblue}{y}-3z=-3 \qquad \rightarrow \qquad 4x-(\color{royalblue}{-2-x-3z})-3z=-3\] Confused about any of that yet? We solved for y in one equation, and replaced that y in another equation with its equivalent in x and z. We now have an equation with 2 variables, not 3.

OpenStudy (anonymous):

(-1, 2, -1) (1, -2, 1) (-1, -2, 1) (1, -2, -1) These are the answers

zepdrix (zepdrix):

lol